OpenStudy (yisrael012332):

can someone help me? Solve for x: −4|x + 5| = −16

OpenStudy (anonymous):

same with\[|x+5| = 4\] since you can divide both side with -4. \[x+5=4 \] \[x+5=-4 \] then just solve those.

OpenStudy (jdoe0001):

\(\bf -4|x + 5| = -16\implies |x + 5| =\cfrac{\cancel{ -16 }}{\cancel{ -4 }} \\ \quad \\ |x + 5| =4\to \begin{cases} +(x+5)=4\\ \bf -(x+5)=4 \end{cases}\)

OpenStudy (anonymous):

 and since its a absolute value you need two answers so make the next one equal to positive 16 and to the same steps